### v1

# Two-Scale Latent Dynamics for Recurrent-Depth Transformers

Abstract
Recurrent-depth transformers scale test-time compute by iterating latent computations before emitting tokens. We study the geometry of these iterates and argue for a simple, two-scale operational picture: (i) within a looped block, updates act as small-scale refinements; (ii) across consecutive blocks, states undergo a larger-scale drift. Across checkpoints, our measurements show that loop steps become smaller and increasingly orthogonal to one another, indicating better local modeling of fine structure rather than merely pushing in a single direction. These dynamics motivate an early-exit mechanism based on the model’s second-order difference in step-size, which we show is superior in terms of performance, stability and time-efficiency, when compared to the KL-divergence exit strategy of Geiping et al. [5] and its naive first-order counterpart.
1 Introduction
How should a language model spend its test-time compute (TTC)? One option is to generate more tokens, externalizing “thinking” into text as in recent chain-of-thought and deliberation systems [9, 1], but this quickly inflates latency and context length. A complementary option is to think in latent space: iterate internal computations before emitting a token so that extra FLOPs improve the representation rather than the sequence length. Recent work shows that looping a Transformer block in this way can yield better reasoning at inference-selected iteration counts [5].
This paper asks a simple question: what geometry emerges when a block is looped? Our thesis is that recurrent depth naturally separates compute across two scales. Within a looped block the model performs local refinements: updates become small and increasingly orthogonal, tracing a stable-curvature spiral around a nearby representation. Across blocks the representation undergoes a slower drift, reflecting coarser changes introduced by depth.
We support this with iterate-only measurements, step norms and consecutive-step angles, tracked over training and visualized with PCA trajectories. Across three recurrent regions in a GPT-2–like model, we consistently observe (i) rapid decay of step size within 551010 loop steps and (ii) stabilization of the angle statistic at moderate values, indicating non-collinear refinements rather than repeated pushes in one direction. PCA plots echo this picture: tight arcs inside loops, larger jumps at block hand-offs.
This paper. We take a geometry-first look at looped blocks. Using only iterate statistics, we show that as training proceeds: (i) loop steps shrink quickly (small-scale refinements), and (ii) consecutive steps become more orthogonal (updates sweep around rather than continue straight). Across blocks, representations change more coarsely (slow drift). Motivated by this pattern, we propose a simple, second-order early-exit rule based on the norm of the difference between consecutive loop updates (“acceleration”). Unlike norms or KL on decoded distributions, acceleration triggers precisely when the local spiral stabilizes, enabling earlier, quality-preserving exits.
Wrapping up, our contributions are three-fold:
  • A diagnostic for loop geometry (step norms and consecutive-step angles) showing shrinking updates and increasing orthogonality over training.
  • Evidence, across three recurrent regions of a GPT-2–like model, that looped blocks perform fine-grained refinements while cross-block changes produce slower drift.
  • A second-order exit criterion that halts when consecutive updates stop changing, delivering a better latency-quality Pareto than step-norm and KL-based baselines.
2 Experimental setup
Backbone and data.
We train a GPT-2-style decoder-only transformer [10, 8] (12 layers, 12 heads, hidden size 768, block size 512) using the same web-text pretraining setup for all experiments. When initializing parts of the model that are applied recurrently (recurrent regions) we use the looped-input projection technique of Geiping et al. [5]. For independently recurring blocks, we inject random noise each time the block is applied (combined with the input sequence) mirroring Geiping et al.’s method for a single repeating block, but extended here to each block group.
Recurrence pattern.
We enable recurrence in three groups: layer 4 (self-loop), layers 5–6 (paired loop), and layer 7 (self-loop). All other layers run single pass.
We collect loop trajectories on held-out text. For tokens and loop step $k$ in group $g$, with hidden $x^{(k)}_{g}\!\in\!\mathbb{R}^{d}$, we define
$\Delta^{(k)}_{g}:=x^{(k+1)}_{g}-x^{(k)}_{g}\,$. (1)
We track (i) step size $\|\Delta^{(k)}_{g}\|_{2}$ and (ii) step orthogonality $\cos\angle\!\big(\Delta^{(k)}_{g},\Delta^{(k-1)}_{g}\big)$. Statistics are averaged across tokens with $1\sigma$ bands.
What we measure.
We focus on two model-agnostic, iterate-only diagnostics: radial refinement, the rapid decay of $\|\Delta^{(k)}\|_{2}$ within a small number of loop steps; and angular refinement, the stabilization of $\cos\angle\!\big(\Delta^{(k)},\Delta^{(k-1)}\big)$ indicating non-collinear, refinement-like updates.
Refer to captionRefer to caption
Figure 1: Overall trajectory with inset zoom (2D PCA). Tight loop refinements (inset) vs. larger cross-block moves (main view).
3 Empirical observations and loop dynamics
We observe three consistent patterns across block groups (4, 5-6, 7) and checkpoints: (i) loop-step sizes shrink rapidly (the mean $\|\Delta^{(k)}\|_{2}$ decays quickly with $k$, typically within 551010 steps), which is consistent with small, diminishing refinements; (ii) consecutive updates become more orthogonal: $\cos\angle\!\big(\Delta^{(k)},\Delta^{(k-1)}\big)$ rises from noisy/low values and settles at lower levels ($\approx 0.5$$0.65$ in later checkpoints), indicating complementary rather than collinear pushes; and (iii) this separation strengthens over training, with faster norm decay and steadier angles in later checkpoints, while across-block changes remain comparatively larger (slow drift).
Supporting visualization: PCA trajectories.
[…]
For each sequence, we refit PCA on the combined set of states so that movements along the loop and depth dimensions are directly comparable. We observe a consistent pattern: tight arcs appear within each loop, while larger jumps mark the transitions between blocks.
(a) Norms Refer to caption Refer to caption Refer to caption
(b) Angles Refer to caption (i) Group 4 Refer to caption (ii) Group 5–6 Refer to caption (iii) Group 7
Figure 2: Loop dynamics. Rows show (a) step norms and (b) step angles. Within each row, panels (i)–(iii) correspond to groups 4, 5–-6, and 7, respectively.
4 Geometry-inspired early exit and trade-offs
We evaluate three exit rules and highlight their computational cost and typical failure modes. In particular:
  1. 1.
    Step-norm (Geiping et al. [5]): trigger when the update magnitude is small,
    $\|\Delta^{(k)}\|_{2}<\tau\,.$ (2)
    Pros: cheap ($\mathcal{O}(d)$ per token-step, where $d$ is hidden size), no decoding. Cons: sensitive to feature scaling; can mis-halt under spiral dynamics where speed plateaus but direction keeps changing. This approach requires a minimum loop count (e.g., $k\!\geq\!2$); an optional normalized variant $\tilde{s}^{(k)}=\|\Delta^{(k)}\|_{2}/(\|x^{(k)}\|_{2}+\varepsilon)$ reduces scale sensitivity.
  2. 2.
    KL-divergence (Geiping et al. [5]): trigger when decoded distributions stabilize,
    $\mathrm{KL}\!\left(p^{(k)}\middle\|p^{(k-1)}\right)<\tau,\quad p^{(k)}=\mathrm{softmax}(z^{(k)})\,.$ (3)
    Pros: tied to the model’s predictive distribution; largely invariant to latent rescalings. Cons: requires decoding over the vocabulary ($\mathcal{O}(V)$); depends on calibration/temperature; reacts later when small latent rotations still change logits. The approach is stable across a wide $\tau$ range but typically slower due to the softmax/KL.
  3. 3.
    Acceleration (ours): detect when the change of the update becomes small,
    $a^{(k)}:=\big\|\Delta^{(k)}-\Delta^{(k-1)}\big\|_{2}\,$, (4)
    and exit when $a^{(k)}<\tau$ for two consecutive steps (i.e., $a^{(k)}<\tau$ and $a^{(k-1)}<\tau$; requires $k\!\geq\!2$). Pros: cheap like step-norm ($\mathcal{O}(d)$); directly sensitive to curvature/rotation. Cons: still affected by global rescaling of latents (mitigated by using the same $\tau$ across blocks or a normalized variant). We use a two-hit check to avoid spurious single-step dips; a bounded, normalized form
    $\hat{a}^{(k)}=\frac{\|\Delta^{(k)}-\Delta^{(k-1)}\|_{2}}{\|\Delta^{(k)}\|_{2}+\|\Delta^{(k-1)}\|_{2}+\varepsilon}\in[0,2]$ (5)
    is a drop-in replacement when per-block scaling differs.
We showed in Section 3 that looped blocks learn to make small, increasingly orthogonal steps, effectively tracing a spiral with nearly constant curvature. A pure norm threshold can be fooled by this regime (norms plateau at a small nonzero value due to rotation), and KL responds later because it measures the decoded distribution. The second-order signal $a^{(k)}$ drops precisely when the local curvature stabilizes; a two-hit check avoids spurious exits without smoothing.
Empirical comparison.
Figure 3 summarizes the trade-offs. We see that (i) acceleration achieves large gains as $\tau$ increases (e.g., $\sim\!580\!\rightarrow\!360$ ms/token from $10^{-5}$ to $10^{-2}$), while KL remains comparatively slow. Further, (ii-iii) acceleration and KL keep PPL/CE essentially flat across thresholds, while step-norm degrades beyond $10^{-3}$ and sharply at $10^{-2}$. These tests show that, at same quality, acceleration dominates KL on latency and avoids the step-norm quality cliff.
(i) Latency vs. threshold.
(ii) Perplexity vs. threshold.
(iii) Cross-entropy vs. threshold.
Figure 3: Exit policies: step-norm (blue), KL (orange), acceleration (green). Acceleration preserves quality while enabling more aggressive thresholds and lower latency; KL preserves quality but is slower; step-norm is fast but loses quality at high $\tau$.
To summarize, the geometry of the loop dynamics naturally motivates a second-order exit criterion. Empirically, this approach achieves a better latency-quality Pareto than KL-based methods and avoids the instability of step-norm under aggressive thresholds. As a result, the method is more robust to threshold variation, reducing or even eliminating the need for extensive hyperparameter tuning.
5 Conclusions and limitations
We provided a geometry-first view of recurrent depth: within looped blocks, updates shrink and stabilize (small, curvature-consistent refinements), while across blocks, representations drift more gradually. Building on this, we introduced a decoding-free, second-order exit that halts when consecutive updates stop accelerating. At matched quality, it lowers latency compared to the KL-based exit of Geiping et al. [5] and avoids the instability of simple step-norm thresholds, while remaining computationally cheap.
Limitations.
Our evidence is observational and iteration-based rather than mechanistic; PCA reduces geometry to 2D; experiments use a single GPT-2-scale model and a specific recurrence pattern; and measurements are averaged over tokens. Despite these caveats, the takeaway is simple and actionable: looped blocks handle local refinements, depth drives slow drift, and curvature stabilization offers a reliable stop signal. Future work includes scaling to larger models and datasets, learning per-block calibration for the exit rule, and integrating geometry-aware controllers with dynamic-depth or token-level TTC strategies.
6 Acknowledgements
This work is partly supported by the MUR FIS2 grant n. FIS-2023-00942 ”NEXUS” (cup B53C25001030001), and by Sapienza University of Rome via the Seed of ERC grant ”MINT.AI” (cup B83C25001040001).
Appendix A Related work
Large language models deliver strong performance but incur high inference cost. One route scales test-time compute in token space by letting models generate longer chains of thought or deliberations, as in OpenAI’s o1 and recent RL-for-reasoning works [9, 1]. This can improve quality but spends computation on extra tokens and quickly saturates the context window.
A complementary direction is to allocate compute in latent space. Classic recurrent computation predates Transformers [7], while within the Transformer family there are two broad threads. (i) Dynamic depth & early exit. Universal Transformers share parameters across depth and learn to halt (often via ACT) [2]; depth-adaptive and early-exit variants similarly modulate per-token compute by stopping when intermediate predictions stabilize (e.g., entropy/confidence heuristics) [3, 11]. (ii) Recurrent depth in latent space. Iterating a recurrent block to “think longer” without emitting tokens improves reasoning while letting inference decide the number of inner steps [5]. Related ideas explore learned latent trajectories and continuous latent-space reasoning [4, 6], and hierarchical controllers that organize multi-stage computation [12].
Appendix B Architecture and training details
Backbone.
We use a GPT-2–style decoder-only transformer [10] implemented in a minimalist NanoGPT setup [8] with learned positional embeddings (no ALiBi/rotary), 12 layers, 12 heads, hidden size d=768d{=}768, and block size 512. Nonlinearities are SiLU. All experiments use the same FineWeb pretraining mixture (identical budgets across runs). We train up to 60 thousand iterations (a total of 300M tokens).
Recurrent regions and loop sampling.
Three depth regions are recurrent: layer 4 (self-loop), layers 5–6 (paired loop, stepped jointly), and layer 7 (self-loop). During training, for each recurrent group $g$ we sample the number of loop iterations $L_{g}$ per forward pass from the lognormal-Poisson schedule defined by Geiping et al. [5], using rate parameter r=12r{=}12. For the paired group (5–6) a single draw governs both layers to preserve coordination. During evaluation, we either fix $L_{g}$ to a budget or use the early-exit criteria in the main text.
Initialization and recurrent input projection.
We adopt the custom initialization for recurrent application described by Geiping et al. [5], and their looped-input projection when re-feeding hidden states to a looping block. For independently recurring blocks we also inject fresh noise when feeding the recurrent input (as in Geiping et al. [5] for a single repeating block), extended here to each block group.
Training recipe.
Unless noted otherwise, optimizer, schedules, and regularization follow our NanoGPT baseline [8]. The only differences are: (i) custom initialization,(ii) SiLU activations, (iii) recurrent loop sampling via lognormal-Poisson (r=12r{=}12), (iv) the recurrent initialization/projection above and (v) sandwhich normalization (using RMSnorm).
Summary. Model: GPT-2–like (12L, 12H, d=768d{=}768), learned PEs, SiLU, FineWeb; recurrence at 4, 5–6, 7; loop counts $L_{g}\sim\text{LogNormalPoisson}(r{=}12)$ per group per pass; recurrent init & projections as in Geiping et al. [5].
Appendix C Additional plots
C.1 Full $\Delta$ angle curves (all checkpoints)
Figure 4 reports the full consecutive-step cosine $\cos\angle\!\big(\Delta^{(k)},\Delta^{(k-1)}\big)$ for groups 4, 5–6, and 7 across all checkpoints.
Figure 4: Full $\cos\angle(\Delta^{(k)},\Delta^{(k-1)})$ curves across checkpoints for groups 4, 5–6, 7. Mean across 10 different samples.
Appendix D Early-exit implementation details
We sweep $\tau\!\in\!\{10^{-5},10^{-4},10^{-3},10^{-2}\}$ (log scale). Latency is average ms/token. Quality is perplexity (PPL) and cross-entropy (CE) on held-out text. The model and recurrence pattern match §2. Each point averages identical prompts across exits.
Algorithm 1 Acceleration-Based Two-Hit Early Exit for Recurrent Depth
0: Block map $f$, initial hidden state $x_{0}$, threshold $\tau>0$, max steps $k_{\max}$
0: Final state $x^{\star}$, steps used $k$
[…]
2:while $k<k_{\max}$ do
3:  $x_{1}\leftarrow f(x_{0})$ {one loop step}
4:  $\delta_{\text{cur}}\leftarrow x_{1}-x_{0}$ {current update}
5:  if $\delta_{\text{prev}}\neq\text{None}$ then
6:   $a\leftarrow\|\delta_{\text{cur}}-\delta_{\text{prev}}\|_{2}$ {acceleration}
7:   $\texttt{small}\leftarrow(a<\tau)$
8:   if small and prev_small then
9:    break {two-hit exit}
10:   end if
11:   $\texttt{prev\_small}\leftarrow\texttt{small}$
12:  end if
13:  $\delta_{\text{prev}}\leftarrow\delta_{\text{cur}}$
14:  $x_{0}\leftarrow x_{1}$
15:  $k\leftarrow k+1$
16:end while
[…]